# IPPP/25/93 Economical Jet Taggers - Equivariant, Slim, and Quantized Antoine Petitjean1, Tilman Plehn1,2, Jonas Spinner3, and Ullrich Köthe2

### v1

@@ Abstract
Modern machine learning is transforming jet tagging at the LHC, but the leading transformer architectures are large, not particularly fast, and training-intensive. We present a slim version of the L-GATr tagger, reduce the number of parameters of jet-tagging transformers, and quantize them. We compare different quantization methods for standard and Lorentz-equivariant transformers and estimate their gains in resource efficiency. We find a six-fold reduction in energy cost for an moderate performance decrease, down to 1000-parameter taggers. This might be a step towards trigger-level jet tagging with small and quantized versions of the leading equivariant transformer architectures.
@@ 1 Introduction
Modern machine learning (ML) is reshaping the research program at the LHC. Even for standard analyses, triggering, data acquisition, object identification, first-principle simulations, and optimal inference are changing rapidly [Plehn:2022ftl]. Among the most established ML applications in LHC physics are jet taggers [Kasieczka:2019dbj, Nachman:2022emq]. Their goal is to utilize the complete information about the jet substructure to determine their partonic nature [Cogan:2014oua, Baldi:2014kfa, deOliveira:2015xxd, Gallicchio:2010sw, Kasieczka:2017nvn, Qu:2019gqs]. This question leads directly to the input and latent data representations, from image-inspired architectures to graphs and transformers for a permutation-invariant point cloud. In addition, jet taggers were shown to benefit from the fundamental Lorentz symmetry [Butter:2017cot]. This motivated a series of Lorentz-equivariant graph-networks taggers [Gong:2022lye, Qiu:2022xvr, Bogatskiy:2022czk, Qiu:2023ihi, ruhe2023clifford], where in physics we would use the term ‘covariant’ rather than ‘equivariant’. The final step of the development of equivariant taggers is marked by Lorentz-equivariant transformers with learned symmetry breaking, using either a geometric algebra representation [Brehmer:2024yqw] or symmetry-linked local reference frames for each constituent [Favaro:2025pgz]. Their advantage over the corresponding graph networks is that transformers benefit significantly more from larger training datasets.
@@
The first and immediate question is how small and how fast we can make an equivariant transformer, assuming that symmetry-aware data representations will improve training and performance under constraints. In Sec. 2, we present and benchmark a streamlined version of L-GATr that uses a more efficient latent representation based only on scalars and vectors. From a representation learning perspective, the slim Lorentz-equivariant tagger can be applied to many LHC tasks, from jet tagging to amplitude regression and event generation. In Sec. 2.5 we compare the performance of L-GATr-slim and LLoCa taggers [Favaro:2025pgz] down to 1000-parameter networks, either with a reduced number of layers or with a reduced numbers of parameters per layer.
The comparison between the L-GATr-slim and LLoCa taggers provides us with a baseline for further computational optimization in Sec. 3. In this second step we quantize the equivariant taggers. Optimized parameter types are an established strategy for large language models, using different precision for different network layers to reduce computational cost and energy consumption. In Sec. 3.2 we introduce weight quantization of the linear layers via a regularization constraint implemented with the proximal gradient method, leading to piecewise-affine regularized quantization (PARQ), and straight-through estimation (STE) as a special case. We apply this quantization to Lorentz-equivariant taggers in Sec. 3.3. We find that they can be quantized like standard networks, leading to a significant efficiency gain with minimal performance loss. Finally, in Sec. 3.4 we combine both strategies to develop truly minimal, i.e. small and quantized versions of the leading tagger architectures. Such maximally resource-efficient versions might become useful for triggering at the HL-LHC.
@@ 2 L-GATr-slim
The standard data representation for neural networks applied to LHC physics is point clouds, including 4-vectors for each particle, jet constituent, or partons. The two symmetries governing these 4-vectors are permutations and Lorentz transformations. Graph-based and transformer architectures ensure permutation invariance. Explicit Lorentz equivariance with learned symmetry breaking in the network architecture has been shown to lead to significant performance improvements for graph networks [Gong:2022lye, Qiu:2022xvr, Bogatskiy:2022czk, Qiu:2023ihi, ruhe2023clifford] and for transformers [Brehmer:2024yqw, Favaro:2025pgz]. The key advantage of transformers over other graph network architectures is their superior scaling: they train efficiently on large datasets thanks to highly parallelizable attention, often yielding the best performance as data and model size grow.
@@ 2.1 Lorentz-scalars and vectors
Finally, we construct the scalar attention closely following L-GATr,
@@ 2.2 Jet tagging
Because our focus will be jet tagging, we first show the L-GATr-slim performance for the standard top tagging benchmark [Kasieczka:2019dbj]. We compare our slim architecture to other leading architectures in Tab. 1 and find that it matches the performance of L-GATr, the LLoCa-Transformer, and different Lorentz-equivariant graph networks. The hyperparameters for the full-sized L-GATr-slim are given in App. A. For details on the established networks and their hyperparameters we refer to Ref. [Favaro:2025pgz]. Because detector effects violate Lorentz symmetry, the tagger has to be able to incorporate the corresponding symmetry breaking. To allow for that, we keep the beam and time axes as an additional input particles.
@@
Next, we benchmark L-GATr-slim on the larger JetClass dataset, covering light quark or gluon jets, , , and jets, as well as top jets. The dataset contains 10M training jets for each of the ten different jet labels. The performance metric is the AUC averaged over all pairs of classes. Again, L-GATr-slim reproduces the performance of other leading Lorentz-equivariant transformers. In addition, L-GATr-slim has the same computational cost as the LLoCa-Transformer and is significantly more efficient than L-GATr, in terms of FLOPs, memory consumption, and training time.
The developments of modern jet taggers from the classic 2017 BDT is illustrated in the left panel of Fig. 1. We note that the highest-performing taggers, denoted as f.t. for fine-tuned, are pretrained on increasingly large datasets, as can be seen in Tab. 1. In the right panel of Fig. 1 we show the same taggers, measuring the performance in AUC for a given training time on a single H100 GPU. This shows that performance is not automatically correlated with computing cost and motivates the development of resource-efficient networks for LHC analyses.
@@ 2.5 Ultra-mini taggers
Finally, we study equivariant transformers for much fewer parameters. This downscaling is motivated by the potential use of transformers at different trigger levels, where resources are scarce [Odagiu:2024bkp]. This is why we go back to jet tagging to ask what fraction of network parameters we can spare at limited computational cost. We train a standard transformer, ParT, the LLoCa-Transformer, and both L-GATr versions with a decreasing number of 2M, 200k, 20k, and 2k parameters, see App. A. For even smaller parameter numbers we refer to Refs. [Bogatskiy:2023fug] and [Vent:2025ddm].
In the left panel of Fig. 3 we see that when the number of network blocks and width are decreased jointly, the LLoCa-Transformer emerges as the leading architecture at small network sizes. L-GATr-slim falls behind the other networks, because of the limited number of operations that exchange information between the scalar and vector channels. Pretraining on JetClass and fine-tuning on the top-tagging dataset yields a clear performance gain for the 1000-parameter LLoCa Transformer, while the 1000-parameter standard transformer sees only a marginal improvement.
@@ 3 Quantized jet taggers
The performance of reduced-size taggers motivates further steps to reduce computational cost, specifically for the LLoCa-Transformer tagger. The bottleneck operation in neural networks is typically the linear layer, which involves large matrix multiplications. So-called quantization performs these matrix multiplications at reduced precision, such that they become faster and require less memory. For other network operations reduced precision would hurt performance, so they should be kept at float16 or even float32 precision. We discuss how to quantize four example transformer-based tagging architectures:
@@ 3.1 Low-precision data types
| float16/bfloat16 | - | 312 | 1513 |
| float8 | - | 624 | 3026 |
@@
The speed of GPU operations for all data types has been increasing continuously, as illustrated in Tab. 4. From the rate of this acceleration we see that modern GPUs are increasingly optimized for smaller data types, such as bfloat16 or float8 instead of the established single-precision float32. For scientific applications, most networks are still trained with single precision, i.e. float32. If we want to reduce the computational cost of neural networks, we can first target the data type used throughout the high-dimensional latent space. This can significantly reduce computational cost, while maintaining performance, as demonstrated for bfloat16 [kalamkar2019study], float8 [micikevicius2022fp8], and float4 [liu2023llm].
@@
During training, we apply the automatic mixed precision strategy of using full precision for specific precision-sensitive operations, as well as the backward pass [micikevicius2017mixed]. We do not reduce the precision in input and output layers to enable the network to process and output high-precision numbers. However, we use float8 precision in the linear layers, which dominate the computational cost. To further speed up training, we could perform all operations at lower precision, but would have to use stochastic rounding [gupta2015deeplearninglimitednumerical]. In general, the push to lower-dimensional latent representations saturates once the network starts to map out the full range of the provided data type. In practice, this happens around float8 precision [wang2023bitnet, ma2024era].
Our implementation emulates low-precision data types in linear layers, because linear layers on float8 are not widely supported yet. We use automatic mixed precision throughout the network and quantize the input of linear layers to the range of float8 values before performing the linear operation in bfloat16 precision. We keep a joint learnable scale parameter in bfloat16 for the final rescaling of the outputs from the float8 range to their original range [wang2023bitnet]. Weights are quantized in the same manner. We use straight-through estimation to propagate gradients through this non-differentiable quantization operation. For our three architectures this means:
@@
Transformer: We first use automatic mixed precision throughout. For all linear layers except the input and output projections, we quantize inputs to float8 using straight-through estimation.
@@
L-GATr-slim: We apply automatic mixed precision and float8 quantization just as for the transformer.
@@
LLoCa-Transformer: The LLoCa-Transformer uses the vanilla transformer as its backbone, and we apply automatic mixed precision and float8 quantization accordingly. The linear layers in the Frames-Net are quantized to float8, whereas the orthonormalization procedure is performed in full float32 precision, to avoid numerical symmetry violation. The frame-to-frame transformations in the attention are also performed in float32 precision.
@@ 3.2 Weight quantization
We compare PARQ and STE for our applications and typically find that PARQ stabilizes training and achieves marginally better performance.
@@ 3.3 Quantized equivariant taggers
First, we compare quantization-aware training with PARQ and with the naive STE in Tab. 5. The vanilla transformer maintains performance up to float8 precision, but drops in performance after weight quantization. In contrast, all networks using Lorentz equivariance maintain constant performance when quantized. Here, it is important to ensure that weight quantization and input quantization violate the Lorentz only equivariance in a minimal way. For the LLoCa-Transformer, we perform the equivariance-critical orthonormalization and the projection into the local frames in full precision, and limit low-precision operations to Lorentz-invariants. For L-GATr-slim, we multiply full vectors with quantized weights, which again does not incur any additional Lorentz violation. Nevertheless, our quantized LLoCa-Transformer and L-GATr-slim can only satisfy Lorentz-equivariance to the precision of the linear layer inputs, which are float8, translating into violation effects at the 10% level. We find that this violation does not translate into significant drops in performance.
@@
| fp32 | fp8 | fp8+STE | fp8+PARQ | |
| Transformer | 1613 118 | 1568 54 | 1186 41 | 1149 55 |
| LLoCa-Transformer | 2150 130 | 2142 136 | 2024 155 | 2049 45 |
| L-GATr-slim | 2264 93 | 2199 62 | 2023 102 | 2145 89 |
For top tagging with resource-efficient networks we estimate the energy use. An H100 has a maximum power consumption of W for the throughput in teraFLOPS given in Tab. 4. We estimate the energy use per operation as for each data type and the total energy consumption of a forward pass by counting all operations in the network and weighting them with the corresponding energy cost [6757323]. For a conservative estimate, we assume that addition and multiplication consume the same amount of energy, effectively reducing the computational cost of linear operations of a float8 latent representation with ternary weights by a factor of two compared to float8-precision weights which also require a multiplication.
Next, we correlate the energy consumption with the performance of different quantized taggers in Fig. 5. Our estimates for the energy cost indicate that every quantization step reduces the cost by a factor of about two, accumulating to an almost order-of-magnitude gain when combining all steps. L-GATr-slim consumes about 50% more energy than a vanilla transformer due to its more expensive linear layers, whereas the LLoCa Transformer incurs a negligible additional energy cost. Importantly, these energy savings come at no performance loss, they only require the implementation of optimized low-precision operations.

### v2

# IPPP/25/93 Economical Jet Taggers – Equivariant, Slim, and Quantized Antoine Petitjean1, Tilman Plehn1,2, Jonas Spinner3, and Ullrich Köthe2

@@ Abstract
Modern machine learning is transforming jet tagging at the LHC, but the leading transformer architectures are large, not particularly fast, and training-intensive. We present a slim version of the L-GATr tagger, reduce the number of parameters of jet-tagging transformers, and quantize them. We compare different quantization methods for standard and Lorentz-equivariant transformers and estimate their gains in resource efficiency. We find an order-of-magnitude reduction in energy cost for an moderate performance decrease, down to 1000-parameter taggers. This might be a step towards trigger-level jet tagging with small and quantized versions of the leading equivariant transformer architectures.
@@ 1 Introduction
Modern machine learning (ML) is reshaping the research program at the LHC. Even for standard analyses, triggering, data acquisition, object identification, first-principle simulations, and optimal inference are changing rapidly [Plehn:2022ftl]. Among the most established ML applications in LHC physics are jet taggers [Kasieczka:2019dbj, Nachman:2022emq]. Their goal is to utilize the complete information about the jet substructure to determine their partonic nature [Cogan:2014oua, Baldi:2014kfa, deOliveira:2015xxd, Gallicchio:2010sw, Kasieczka:2017nvn, Qu:2019gqs]. This question leads directly to the input and latent data representations, from image-inspired architectures to graph networks and transformers for permutation-invariant point clouds. In addition, jet taggers were shown to benefit from the fundamental Lorentz symmetry [Butter:2017cot]. This motivated a series of Lorentz-equivariant graph-networks taggers [Gong:2022lye, Qiu:2022xvr, Bogatskiy:2022czk, Qiu:2023ihi, ruhe2023clifford], where in physics we would use the term ‘covariant’ rather than ‘equivariant’. The final step of the development of equivariant taggers is marked by Lorentz-equivariant transformers with learned symmetry breaking, using either a geometric algebra representation [Brehmer:2024yqw] or symmetry-linked local reference frames for each constituent [Favaro:2025pgz]. Their advantage over the corresponding graph networks is that transformers benefit significantly more from larger training datasets and scale more efficiently with network size.
@@
The first and immediate question is how small and how fast we can make an equivariant transformer, assuming that symmetry-aware data representations will improve training and performance under constraints. In Sec. 2, we present and benchmark a streamlined version of the L-GATr architecture that uses a more efficient latent representation based only on scalars and vectors. This slim Lorentz-equivariant tagger can be applied to many LHC tasks, from jet tagging to amplitude regression and event generation. In Sec. 2.5 we compare the performance of L-GATr-slim and LLoCa taggers [Favaro:2025pgz] down to 1000-parameter networks, either with a reduced number of layers or with a reduced numbers of parameters per layer.
The comparison between the L-GATr-slim and LLoCa taggers provides us with a baseline for further computational optimization in Sec. 3. In this second step we quantize the equivariant taggers. Optimized data types are an established strategy for large language models, using different precision for different network layers to minimize computational cost and energy consumption. In Sec. 3.2 we introduce weight quantization of the linear layers via a regularization constraint implemented with the proximal gradient method, leading to piecewise-affine regularized quantization (PARQ), and straight-through estimation (STE) as a special case. We apply this quantization to Lorentz-equivariant taggers in Sec. 3.3. We find that they can be quantized like standard networks, leading to a significant efficiency gain with minimal performance loss. Finally, in Sec. 3.4 we combine both strategies to develop truly minimal, i.e. small and quantized versions of the leading tagger architectures. Such maximally resource-efficient versions might become useful for triggering at the HL-LHC.
@@ 2 L-GATr-slim
The standard data representation for neural networks applied to LHC physics is point clouds, including 4-vectors for each particle, jet constituent, or partons. The two symmetries governing these 4-vectors are permutations and Lorentz transformations. Graph-based and transformer architectures ensure permutation invariance. Explicit Lorentz equivariance with learned symmetry breaking in the network architecture has been shown to lead to significant performance improvements for graph networks [Gong:2022lye, Qiu:2022xvr, Bogatskiy:2022czk, Qiu:2023ihi, ruhe2023clifford] and for transformers [Brehmer:2024yqw, Favaro:2025pgz]. The key advantage of transformers over other graph network architectures is their superior scaling: they train efficiently on large datasets thanks to highly parallelizable attention, often yielding the best performance as data and network size grow.
@@ 2.1 Lorentz-scalars and vectors
Finally, we construct the scalar attention matrix closely following L-GATr,
@@ 2.2 Jet tagging
Because our focus will be jet tagging, we first show the L-GATr-slim performance for the standard top tagging benchmark [Kasieczka:2019dbj]. We compare our slim architecture to other leading architectures in Tab. 1 and find that it matches the performance of L-GATr, the LLoCa-Transformer, and different Lorentz-equivariant graph networks. The hyperparameters for the full-sized L-GATr-slim are given in App. A. For details on the established networks and their hyperparameters we refer to Ref. [Favaro:2025pgz]. Because detector effects violate Lorentz symmetry, the tagger has to be able to incorporate the corresponding symmetry breaking. To allow for that, we keep the beam and time axes as additional input particles [Brehmer:2024yqw, Favaro:2025pgz].
@@
| Rej50% | Rej50% | Rej50% | Rej50% | Rej99% | Rej50% | Rej99.5% | Rej50% | Rej50% | |
| MIParT-L [Wu:2024thh] | 10753 | 4202 | 123 | 1927 | 5450 | 31250 | 16807 | 542 | 402 |
| LorentzNet* [Favaro:2025pgz] | 8475 | 2729 | 111 | 1152 | 3515 | 13889 | 10257 | 400 | 303 |
| PELICAN-lite* [Favaro:2025pgz] | 8333 | 3040 | 113 | 1321 | 3802 | 17084 | 10363 | 435 | 332 |
| ParticleNet [Qu:2019gqs] | 7634 | 2475 | 104 | 954 | 3339 | 10526 | 11173 | 347 | 283 |
| LLoCa-ParticleNet* [Favaro:2025pgz] | 7463 | 2833 | 105 | 1072 | 3155 | 10753 | 9302 | 403 | 306 |
| ParT [Qu:2022mxj] | 10638 | 4149 | 123 | 1864 | 5479 | 32787 | 15873 | 543 | 402 |
| LLoCa-ParT* [Favaro:2025pgz] | 11561 | 4640 | 125 | 2037 | 5900 | 41667 | 19231 | 552 | 419 |
| Transformer [Favaro:2025pgz] | 10753 | 3333 | 116 | 1369 | 4630 | 24390 | 17857 | 415 | 334 |
| LLoCa-Transformer* [Favaro:2025pgz] | 11628 | 4651 | 125 | 2037 | 5618 | 39216 | 17241 | 548 | 410 |
| L-GATr* [Brehmer:2024yqw] | 12987 | 4819 | 128 | 2311 | 6116 | 47619 | 20408 | 588 | 432 |
| L-GATr-slim* | 12121 | 4843 | 127 | 2309 | 5952 | 40113 | 21053 | 585 | 429 |
Next, we benchmark L-GATr-slim on the larger JetClass dataset, covering light quark or gluon jets, , , and jets, as well as top jets. The dataset contains 10M training jets for each of the ten different jet labels. We report the AUC averaged over all pairs of classes, accuracy, rejection rates as well as computational cost metrics in Tables 2 and 3. Again, L-GATr-slim reproduces the performance of other leading Lorentz-equivariant transformers. In addition, L-GATr-slim has the same computational cost as the LLoCa-Transformer and is significantly more efficient than L-GATr, in terms of FLOPs, memory consumption, and training time.
The development of modern jet taggers from the classic 2017 BDT is illustrated in the left panel of Fig. 1. We note that the highest-performing taggers, denoted as f.t. for fine-tuned, are pretrained on increasingly large datasets, as can be seen in Tab. 1. In the right panel of Fig. 1 we show the same taggers, measuring the performance in AUC for a given training time on a single H100 GPU. This shows that performance is not automatically correlated with computing cost and motivates the development of resource-efficient networks for LHC analyses.
@@ 2.5 Ultra-mini taggers
Finally, we study equivariant transformers for much fewer parameters. This downscaling is motivated by the potential use of transformers at different trigger levels, where resources are scarce [Odagiu:2024bkp]. This is why we go back to jet tagging to ask what fraction of network parameters we can spare at limited computational cost. We train a standard transformer, ParT, the LLoCa-Transformer, and L-GATr-slim with a decreasing number of 2M, 200k, 20k, and 2k parameters, see App. A. For even smaller parameter numbers we refer to Refs. [Bogatskiy:2023fug] and [Vent:2025ddm].
In the left panel of Fig. 3 we see that when the number of network blocks and width are decreased jointly, the LLoCa-Transformer emerges as the leading architecture at small network sizes. L-GATr-slim falls behind the other networks, because of the limited number of operations that exchange information between the scalar and vector channels. Pretraining on JetClass and fine-tuning on the top-tagging dataset yields a clear performance gain for the 1000-parameter LLoCa-Transformer due to the increased expressivity, while the 1000-parameter standard transformer sees only a marginal improvement.
@@ 3 Quantized jet taggers
The performance of reduced-size taggers motivates further steps to reduce computational cost, specifically for the LLoCa-Transformer tagger. The bottleneck operation in neural networks is typically the linear layer, which involves large matrix multiplications. So-called quantization performs these matrix multiplications at reduced precision, such that they become faster and require less memory. For other network operations reduced precision would hurt performance, so they should be kept at bfloat16 or even float32 precision.
Quantized jet taggers have been studied previously, for example in P-DAT-Bit [Krause:2025qnl] and BitParT [Rai:2025cog]. Here we describe how to quantize four representative transformer-based tagging architectures:
@@
ParT: Transformer using edge features in attention, and class attention to aggregate over particles;
@@ 3.1 Low-precision data types
| float16/bfloat16 | – | 312 | 1513 |
| float8 | – | 624 | 3026 |
| [pJ] | [pJ] | |
| float32 | 0.38 | 1.31 |
| bfloat16 | 0.11 | 0.21 |
| int8 | 0.007 | 0.07 |
@@
The speed of GPU operations for all data types has been increasing continuously, as illustrated in Tab. 5. From the rate of this acceleration we see that modern GPUs are increasingly optimized for smaller data types, such as bfloat16, float8 and int8 instead of the established single-precision float32. For scientific applications, most networks are still trained with float32 precision. If we want to reduce the computational cost of neural networks, we can first target the data type used throughout the high-dimensional latent space. This can significantly reduce computational cost, while maintaining performance, as demonstrated for bfloat16 [kalamkar2019study], int8 [jacob2018quantization], float8 [micikevicius2022fp8], and float4 [liu2023llm].
@@
During training, we apply the automatic mixed precision strategy of using full precision for specific precision-sensitive operations, as well as the backward pass [micikevicius2017mixed]. We do not reduce the precision in input and output layers beyond bfloat16 to enable the network to embed and output high-precision numbers. However, we use int8 precision in the hidden linear layers, which dominate the computational cost. To further speed up training, we could also perform the backward pass at lower precision, but would have to use stochastic rounding [gupta2015deep]. In general, the push to lower-dimensional latent representations saturates once the network starts to map out the full range of the provided data type. In practice, this happens around int8 or float8 precision [wang2023bitnet, ma2024era].
Our implementation emulates low-precision data types in linear layers, because linear layers operating on int8 are not widely supported yet. We use automatic mixed precision throughout the network and quantize the input of linear layers to the range of int8 values before performing the linear operation in bfloat16 precision. To quantize the inputs with minimal loss of information, we use zero-point quantization [jacob2018quantization]:
The scale maps the range to ( for int8) and the zeropoint centers the range to align the zero values. denotes rounding and applies the min and max functions with the quantization range. The values and can be computed on the fly for each tensor (dynamic quantization) or pre-computed from training data and fixed at inference (static quantization). We only use static quantization for the small taggers of Section 3.4, as the cost of dynamic quantization is negligible for large networks. We find similar performance when replacing int8 by float8, but prefer int8 because of the slightly reduced computational cost. Weights are quantized in the same manner. We use straight-through estimation to propagate gradients through this non-differentiable quantization operation. For our four architectures this means:
@@
Transformer: We first use automatic mixed precision throughout. For all linear layers except the input and output projections, we quantize inputs to int8 using straight-through estimation.
ParT: We again use automatic mixed precision on the whole architecture. Linear layers in the particle attention and class attention blocks are quantized to int8 in the same manner as above. The feed-forward networks embedding particle-wise and pair-wise features are also quantized to int8, except their first layer which uses bfloat16 precision.
@@
L-GATr-slim: We apply automatic mixed precision and int8 quantization just as for the transformer.
@@
LLoCa-Transformer: The LLoCa-Transformer uses the vanilla transformer as its backbone, and we apply automatic mixed precision and int8 quantization accordingly. The linear layers in the Frames-Net are quantized to int8, whereas the orthonormalization procedure is performed in full float32 precision, to avoid numerical symmetry violation. The frame-to-frame transformations in attention are also performed in float32 precision.
@@ 3.2 Weight quantization
We compare PARQ and STE for our applications and find that PARQ stabilizes training and achieves marginally better performance.
@@ 3.3 Quantized equivariant taggers
First, we compare quantization-aware training with PARQ and with the naive STE in Tab. 6. The vanilla transformer maintains performance up to int8 precision, but drops in performance after weight quantization. ParT maintains performance all the way down to ternary weights. The two Lorentz-equivariant networks undergo slight performance drops when reducing precision to int8, and when quantizing weights to ternary values. Here, it is important to ensure that weight quantization and input quantization violate the Lorentz equivariance in a minimal way. For the LLoCa-Transformer, we perform the equivariance-critical orthonormalization and the projection into the local frames in full precision, and limit low-precision operations to Lorentz-invariants. For L-GATr-slim, we multiply full vectors with quantized weights, which again does not incur any additional Lorentz violation. Nevertheless, our quantized LLoCa-Transformer and L-GATr-slim can only satisfy Lorentz-equivariance to the precision of the linear layer inputs, which are float8, translating into violation effects at the 10% level.
@@
| f32 | i8 | i8+STE | i8+PARQ | |
| Transformer | 1694 69 | 1577 68 | 1357 69 | 1302 55 |
| ParT | 1808 33 | 1768 81 | 1676 94 | 1693 86 |
| LLoCa-Transformer | 2150 130 | 2019 79 | 1932 132 | 1935 110 |
| L-GATr-slim | 2264 93 | 2082 124 | 1990 108 | 1872 67 |
For top tagging with resource-efficient networks we estimate the energy use following Refs. [6757323, wang2023bitnet]. We use estimates for the energy cost per operation in Table 5, and estimate the total energy consumption of a forward pass by counting all operations in the network and weighting them with the corresponding energy cost. We quantize inner linear layers to ternary weights, and reduce the precision of the latent representation as described in Sec. 3.1.
Next, we correlate the energy consumption with the performance of different quantized taggers in Fig. 5. Our estimates for the energy cost indicate that every quantization step reduces the cost by a factor of about two, accumulating to an order-of-magnitude gain when combining all steps. L-GATr-slim consumes about 50% more energy than a vanilla transformer due to its more expensive linear layers, whereas the LLoCa-Transformer incurs a negligible additional energy cost. Compared with the vanilla Transformer, ParT incurs about 50% overhead with roughly equal contributions from preprocessing the pairwise features and from the larger MLP hidden dimension in the Transformer blocks. Importantly, these energy savings come at almost no performance loss, they only require the implementation of optimized low-precision operations.